OpenStudy (anonymous):

sqrt 3x-4=3 a.)7/3 b.)-1/3 c.)13/3

OpenStudy (anonymous):

If I'm reading this right, \[\sqrt{3x-4}=3\] In which case, square both sides \[3x-4=9\] and solve for x \[x=13/3\] So the correct answer would be c.

OpenStudy (anonymous):

Thanks! :-)